in a market for figs (Q, measured in kilograms) monthly demand and supply is given by:
QP (p) = 280,000 – 20,000p QS(p) = 5,000p – 20,000
market equilibrium price is p*= 12
market equilibrium quantity is q* = 40,000

a) compute the price elasticity of supply of figs and the price elasticity of demand of figs at the equilibrium point.
b) do producers or consumers have the relatively less elastic curve in this market?
Answer #1

Qd = 280,000 - 29,000p dQd = 0 -20,000(1. = -26,000 Equilibrium price = 12 Q = 40000 Ed = dad x ap = -260,600 x 12 yoooo Ed =Ed= - 18 २. व SS - - 18,११६ 2 = – ५.२६ Price Elcesticity of demand Ed = -५.२